# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ H.1 Proof of Technical Lemmas
$\left(\bm{Q}^{\star}_{t}-\bm{Q}_{t}^{(A_{t})}\right)^{\top}\bm{\theta}_{Q}$ $=2\sum_{i=1}^{d_{\texttt{lin}}}\mathbbm{1}\{\bm{Q}_{t}=\bm{e}_{i}\}\mathbbm{1}\{A_{t}\neq\mathop{\mathrm{sign}}\left(\bm{\theta}_{Q,i}\right)\}|\bm{\theta}_{Q,i}|$
$=2\sqrt{\frac{d_{\texttt{lin}}}{T}}\sum_{i=1}^{d_{\texttt{lin}}}\mathbbm{1}\{\bm{Q}_{t}=\bm{e}_{i}\}\mathbbm{1}\{A_{t}\neq\mathop{\mathrm{sign}}\left(\bm{\theta}_{Q,i}\right)\}$
where the last equality follows from the fact that $|\theta_{Q,i}|=\sqrt{d_{\texttt{lin}}/T}$ as given in Equation (G.7). Recall the definition of $\mathcal{R}_{t}^{\texttt{lin}}$ in Equation (H.6). Combined with the above display, it follows that
$\mathcal{R}^{\texttt{lin}}_{t}(\bm{\theta},f)$ $=2\sqrt{\frac{d_{\texttt{lin}}}{T}}\sum_{\tau=1}^{t}\sum_{i=1}^{d_{\texttt{lin}}}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{\tau}\neq\mathop{\mathrm{sign}}(\bm{\theta}_{Q,i}),V_{\tau}=1,\bm{Q}_{\tau}=\bm{e}_{i}\right\}\mid\mathcal{H}_{\tau-1}\right]$ (H.17)
$=\sqrt{\frac{1}{d_{\texttt{lin}}T}}\sum_{i=1}^{d_{\texttt{lin}}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{\tau}\neq\mathop{\mathrm{sign}}(\bm{\theta}_{Q,i})\right)\mid\mathcal{H}_{\tau-1},\bm{Q}_{\tau}=\bm{e}_{i}\right]$
@@
as specified by the data generating process in Definition G.1 and Equation (H.4). Consider $\bm{\theta}^{\prime}_{Q}\in{{\mathbb{R}}}^{d_{\texttt{lin}}}$ such that $\theta^{\prime}_{Q,j}=\theta_{Q,j}$ for all $j\neq i$ and $\theta^{\prime}_{Q,i}=-\theta_{Q,i}$. Let $\mathbb{P}^{(t-1)}_{Q,i}:=\mathbb{P}\left(\cdot\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right)$. Continuing from Equation (H.17), by the Bretagnolle-Huber inequality as stated in Theorem H.1, we have for any $t\in[T]$,
$\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]+\mathbb{E}^{(t-1)}_{\bm{\theta}^{\prime},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta^{\prime}_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]$
@@
where the last inequality follows from Equations (H.16), (H.1) and (H.2). Since $\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}\left(f,f\right)=0$, it follows from the above display that
$\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]+\mathbb{E}^{(t-1)}_{\bm{\theta}^{\prime},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta^{\prime}_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]$ (H.18)
@@
where the last inequality follows from Equation (H.5). Let $\Theta_{d_{\texttt{lin}}}\subset{{\mathbb{R}}}^{d_{\texttt{lin}}}$ denote the set of all vectors whose coordinate are either $\beta:=\sqrt{d_{\texttt{lin}}/T}$ or $\beta$, i.e.,
@@
For any vector $\bm{\theta}\in{{\mathbb{R}}}^{d}$ and $j\in[d]$, denote $(\theta_{1},\cdots,\theta_{j-1},\theta_{j+1},\cdots,\theta_{d})\in{{\mathbb{R}}}^{d-1}$ as $\bm{\theta}_{[-j]}$ and $\bm{\theta}^{i}_{[-j]}:=(\theta_{1},\cdots,\theta_{j-1},i,\theta_{1},\cdots,\theta_{d})\in{{\mathbb{R}}}^{d}$ for $i\in{{\mathbb{R}}}$. Applying an average hammer over all $\bm{\theta}_{Q}\in\Theta_{d_{\texttt{lin}}}$, which satisfies $|\Theta_{d_{\texttt{lin}}}|=2^{d_{\texttt{lin}}}$, it follows from Equation (H.17) that
$\sup_{\bm{\theta}\in\Theta}\mathcal{R}^{\texttt{lin}}_{t}(\bm{\theta},f)$ $\geq\frac{1}{|\Theta_{d_{\texttt{lin}}}|}\sum_{\bm{\theta}_{Q}\in\Theta_{d_{\texttt{lin}}}}\sqrt{\frac{1}{d_{\texttt{lin}}T}}\sum_{i=1}^{d_{\texttt{lin}}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{\tau}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{\tau-1},\bm{Q}_{\tau}=\bm{e}_{i}\right]$ (H.19)
$\geq\frac{1}{2^{d_{\texttt{lin}}}}\sum_{i=1}^{d_{\texttt{lin}}}\sum_{\bm{\theta}^{j}_{Q,[-i]}\in\Theta_{d_{\texttt{lin}}}}\sum_{j\in\{\pm\beta\}}\sqrt{\frac{1}{d_{\texttt{lin}}T}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta}^{j}_{Q,[-i]},f,\pi}\mathbb{E}\left[\mathbbm{1}\left(A_{\tau}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{\tau-1},\bm{Q}_{\tau}=\bm{e}_{i}\right]$
@@
where inequality $(i)$ follows from Equation (H.18) and equality $(ii)$ follows from
@@
Taking $t=T$ in Equation (H.19) yields the result in Equation (H.7). ∎
H.1.4 Proof of Lemma H.4
@@
$\left|x^{\beta}-y^{\beta}\right|\leq\beta\left|x-y\right|.$ (H.20)
For $\mathbf{o},\mathbf{o}^{\prime}\in B_{k}$, by definition of $f_{\bm{\omega}}$ in Equation (H.10), we have
@@
where the second equality follows from the definition of $\varphi_{k}$ in Equation (H.11) and the last equality follows from the definition of $\phi_{\beta}$ in Equation (H.8). Continuing from the above display,
$\left|f_{\bm{\omega}}(\mathbf{o})-f_{\bm{\omega}}(\mathbf{o}^{\prime})\right|$ $=2^{\beta}C_{\phi}\left|\left(\frac{1}{2M}-\|\mathbf{o}-\bm{b}_{k}\|_{\infty}\right)^{\beta}-\left(\frac{1}{2M}-\|\mathbf{o}-\bm{b}_{k}\|_{\infty}\right)^{\beta}\right|$ (H.21)
@@
where equality $(i)$ follows from Equation (H.20) and inequality $(ii)$ follows from the triangle inequality.
@@
$\left|f(\mathbf{o})-f(\mathbf{o}^{\prime})\right|$ $\leq\max\left\{\left|f(\mathbf{o})-f(\mathbf{p}_{k})\right|,\left|f(\mathbf{o}^{\prime})-f(\mathbf{p}_{k^{\prime}})\right|\right\}$ (H.22)
@@
where the last inequality follows from Equation (H.21). We can pick $p_{k}$ and $p_{k^{\prime}}$ so that
@@
it then follows from Equation (H.22) that
@@
Combining the above display with Equation (H.21) finishes the proof. ∎
H.1.5 Proof of Lemma H.5
@@
For any $\mathbf{o}\in\widetilde{B}_{j}$, it follows from Equation (H.10) that
$f_{\bm{\omega}}(\mathbf{o})=\omega_{j}\varphi_{j}(\mathbf{o})\geq C_{\phi}M^{-\beta}\delta M^{\beta}=\delta C_{\phi}.$ (H.23)
For any $\bm{\omega}\in\Omega_{m}$ and $\delta>0$, combining Equation (H.23) with $\mathcal{R}_{T}^{\texttt{non}}$ as defined in Equation (H.6) yields that
$\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f_{\bm{\omega}})$ $=\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq\mathop{\mathrm{sign}}\left(f_{\bm{\omega}}(\bm{O}_{t})\right),V_{t}=0\right\}\left|f_{\bm{\omega}}(\bm{O}_{t})\right|\mid\mathcal{H}_{t-1}\right]$ (H.24)
$=\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq\mathop{\mathrm{sign}}\left(f_{\bm{\omega}}(\bm{O}_{t})\right),V_{t}=0\right\}\mathbbm{1}\left\{\bm{O}_{t}\in B_{j}\right\}\left|f_{\bm{\omega}}(\bm{O}_{t})\right|\mid\mathcal{H}_{t-1}\right]$
$\geq C_{\phi}\delta\sum_{j=1}^{m}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq\omega_{j},\bm{O}_{t}\in\widetilde{B}_{j},V_{t}=0\right\}\mid\mathcal{H}_{t-1}\right].$
@@
$\mathbb{P}\left(\bm{O}\in\widetilde{B}_{1}\right)$ $=\mathbb{P}\left(\phi_{\beta}\left(2M[\bm{O}-\bm{b}_{1}]\right)\geq\delta M^{\beta},\bm{O}\in B_{1}\right)$ (H.25)
$=\int_{B_{1}}\mathbbm{1}\left(\phi_{\beta}(2M(\mathbf{o}-\bm{b}_{1}))\geq\delta M^{\beta}\right)d\mathbf{o}$
$=\int_{B_{1}}\mathbbm{1}\left\{(1-2M\|\mathbf{o}-\bm{b}_{1}\|_{\infty})^{\beta}\geq\delta M^{\beta}\right\}d\mathbf{o}$
$=\int_{\left[0,\frac{1}{M}\right]^{d_{\texttt{non}}}}\mathbbm{1}\left(\max_{l\in[d_{\texttt{non}}]}\left|o_{l}-\frac{1}{2M}\right|\leq\frac{1}{2M}-\frac{1}{2}\delta^{1/\beta}\right)d\mathbf{o}$
$=\int_{\left[0,\frac{1}{M}\right]^{d_{\texttt{non}}}}\mathbbm{1}\left(\mathbf{o}\in\left[\frac{1}{2}\delta^{1/\beta},\frac{1}{M}-\frac{1}{2}\delta^{1/\beta}\right]^{d_{\texttt{non}}}\right)d\mathbf{o}$
@@
To handle $\mathcal{K}^{(\texttt{non})}$ as defined in Equation (H.1), take $\bm{\omega}$ and $\bm{\omega}^{\prime}$ so that they only differ in $\omega_{j}$, we have
@@
$\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[(f_{\bm{\omega}}(\bm{O}_{t})-f_{\bm{\omega}^{\prime}}(\bm{O}_{t}))^{2}\mathbbm{1}\left\{A_{t}=1,V_{t}=0\right\}\mid\mathcal{H}_{t-1}\right]$ (H.26)
$~~=4\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\varphi_{j}(\bm{O}_{t})^{2}\mathbbm{1}\left\{A_{t}=1,\bm{O}_{t}\in B_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
$~~\leq\frac{4C^{2}_{\phi}\delta^{2}}{M^{d_{\texttt{non}}}}+4\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\varphi_{j}(\bm{O}_{t})^{2}\mathbbm{1}\left\{A_{t}=1,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
$~~\leq\frac{4C^{2}_{\phi}\delta^{2}}{M^{d_{\texttt{non}}}}+4C_{\phi}^{2}M^{-2\beta}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}=1,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
where the first equality follows from the fact that $\bm{O}_{t}\in B_{j}$ already implies $V_{t}=0$. Similarly, for any $i\in[N]$, applying the above argument with Equation (H.25) to the pretrained data yields
$\mathbb{E}_{f_{\bm{\omega}}}\left[\left(f_{\bm{\omega}}\left(\bm{O}_{i}^{(0)}\right)-f_{\bm{\omega}^{\prime}}\left(\bm{O}_{i}^{(0)}\right)\right)^{2}\right]$ $\leq\frac{4C^{2}_{\phi}\delta^{2}}{M^{d_{\texttt{non}}}}+4C^{2}_{\phi}M^{-2\beta}\left(\frac{1}{M}-\delta^{1/\beta}\right)^{d_{\texttt{non}}}.$ (H.27)

### v3

@@ I.2 Theoretical Guarantees and Regret Comparison
$\left(\bm{Q}^{\star}_{t}-\bm{Q}_{t}^{(A_{t})}\right)^{\top}\bm{\theta}_{Q}$ $=2\sum_{i=1}^{d_{\texttt{lin}}}\mathds{1}\{\bm{Q}_{t}=\bm{e}_{i}\}\mathds{1}\{A_{t}\neq\mathop{\mathrm{sign}}\left(\bm{\theta}_{Q,i}\right)\}|\bm{\theta}_{Q,i}|$
$=2\sqrt{\frac{d_{\texttt{lin}}}{T}}\sum_{i=1}^{d_{\texttt{lin}}}\mathds{1}\{\bm{Q}_{t}=\bm{e}_{i}\}\mathds{1}\{A_{t}\neq\mathop{\mathrm{sign}}\left(\bm{\theta}_{Q,i}\right)\}$
where the last equality follows from the fact that $|\theta_{Q,i}|=\sqrt{d_{\texttt{lin}}/T}$ as given in Equation (J.7). Recall the definition of $\mathcal{R}_{t}^{\texttt{lin}}$ in Equation (K.6). Combined with the above display, it follows that
$\mathcal{R}^{\texttt{lin}}_{t}(\bm{\theta},f)$ $=2\sqrt{\frac{d_{\texttt{lin}}}{T}}\sum_{\tau=1}^{t}\sum_{i=1}^{d_{\texttt{lin}}}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left\{A_{\tau}\neq\mathop{\mathrm{sign}}(\bm{\theta}_{Q,i}),V_{\tau}=1,\bm{Q}_{\tau}=\bm{e}_{i}\right\}\mid\mathcal{H}_{\tau-1}\right]$ (K.17)
$=\sqrt{\frac{1}{d_{\texttt{lin}}T}}\sum_{i=1}^{d_{\texttt{lin}}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{\tau}\neq\mathop{\mathrm{sign}}(\bm{\theta}_{Q,i})\right)\mid\mathcal{H}_{\tau-1},\bm{Q}_{\tau}=\bm{e}_{i}\right]$
@@
as specified by the data generating process in Definition J.1 and Equation (K.4). Consider $\bm{\theta}^{\prime}_{Q}\in{{\mathbb{R}}}^{d_{\texttt{lin}}}$ such that $\theta^{\prime}_{Q,j}=\theta_{Q,j}$ for all $j\neq i$ and $\theta^{\prime}_{Q,i}=-\theta_{Q,i}$. Let $\mathbb{P}^{(t-1)}_{Q,i}:=\mathbb{P}\left(\cdot\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right)$. Continuing from Equation (K.17), by the Bretagnolle-Huber inequality as stated in Theorem K.1, we have for any $t\in[T]$,
$\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]+\mathbb{E}^{(t-1)}_{\bm{\theta}^{\prime},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta^{\prime}_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]$
@@
where the last inequality follows from Equations (K.16), (K.1) and (K.2). Since $\mathcal{K}^{(\texttt{non})}_{\bm{\theta},t}\left(f,f\right)=0$, it follows from the above display that
$\mathbb{E}^{(t-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]+\mathbb{E}^{(t-1)}_{\bm{\theta}^{\prime},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{t}\neq\mathop{\mathrm{sign}}(\theta^{\prime}_{Q,i})\right)\mid\mathcal{H}_{t-1},\bm{Q}_{t}=\bm{e}_{i}\right]$ (K.18)
@@
where the last inequality follows from Equation (K.5). Let $\Theta_{d_{\texttt{lin}}}\subset{{\mathbb{R}}}^{d_{\texttt{lin}}}$ denote the set of all vectors whose coordinate are either $\beta:=\sqrt{d_{\texttt{lin}}/T}$ or $\beta$, i.e.,
@@
For any vector $\bm{\theta}\in{{\mathbb{R}}}^{d}$ and $j\in[d]$, denote $(\theta_{1},\cdots,\theta_{j-1},\theta_{j+1},\cdots,\theta_{d})\in{{\mathbb{R}}}^{d-1}$ as $\bm{\theta}_{[-j]}$ and $\bm{\theta}^{i}_{[-j]}:=(\theta_{1},\cdots,\theta_{j-1},i,\theta_{1},\cdots,\theta_{d})\in{{\mathbb{R}}}^{d}$ for $i\in{{\mathbb{R}}}$. Applying an average hammer over all $\bm{\theta}_{Q}\in\Theta_{d_{\texttt{lin}}}$, which satisfies $|\Theta_{d_{\texttt{lin}}}|=2^{d_{\texttt{lin}}}$, it follows from Equation (K.17) that
$\sup_{\bm{\theta}\in\Theta}\mathcal{R}^{\texttt{lin}}_{t}(\bm{\theta},f)$ $\geq\frac{1}{|\Theta_{d_{\texttt{lin}}}|}\sum_{\bm{\theta}_{Q}\in\Theta_{d_{\texttt{lin}}}}\sqrt{\frac{1}{d_{\texttt{lin}}T}}\sum_{i=1}^{d_{\texttt{lin}}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{\tau}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{\tau-1},\bm{Q}_{\tau}=\bm{e}_{i}\right]$ (K.19)
$\geq\frac{1}{2^{d_{\texttt{lin}}}}\sum_{i=1}^{d_{\texttt{lin}}}\sum_{\bm{\theta}^{j}_{Q,[-i]}\in\Theta_{d_{\texttt{lin}}}}\sum_{j\in\{\pm\beta\}}\sqrt{\frac{1}{d_{\texttt{lin}}T}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta}^{j}_{Q,[-i]},f,\pi}\mathbb{E}\left[\mathds{1}\left(A_{\tau}\neq\mathop{\mathrm{sign}}(\theta_{Q,i})\right)\mid\mathcal{H}_{\tau-1},\bm{Q}_{\tau}=\bm{e}_{i}\right]$
@@
where inequality $(i)$ follows from Equation (K.18) and equality $(ii)$ follows from
@@
Taking $t=T$ in Equation (K.19) yields the result in Equation (K.7). ∎
K.1.4 Proof of Lemma K.4
@@
$\left|x^{\beta}-y^{\beta}\right|\leq\beta\left|x-y\right|.$ (K.20)
For $\mathbf{o},\mathbf{o}^{\prime}\in B_{k}$, by definition of $f_{\bm{\omega}}$ in Equation (K.10), we have
@@
where the second equality follows from the definition of $\varphi_{k}$ in Equation (K.11) and the last equality follows from the definition of $\phi_{\beta}$ in Equation (K.8). Continuing from the above display,
$\left|f_{\bm{\omega}}(\mathbf{o})-f_{\bm{\omega}}(\mathbf{o}^{\prime})\right|$ $=2^{\beta}C_{\phi}\left|\left(\frac{1}{2M}-\|\mathbf{o}-\bm{b}_{k}\|_{\infty}\right)^{\beta}-\left(\frac{1}{2M}-\|\mathbf{o}-\bm{b}_{k}\|_{\infty}\right)^{\beta}\right|$ (K.21)
@@
where equality $(i)$ follows from Equation (K.20) and inequality $(ii)$ follows from the triangle inequality.
@@
$\left|f(\mathbf{o})-f(\mathbf{o}^{\prime})\right|$ $\leq\max\left\{\left|f(\mathbf{o})-f(\mathbf{p}_{k})\right|,\left|f(\mathbf{o}^{\prime})-f(\mathbf{p}_{k^{\prime}})\right|\right\}$ (K.22)
@@
where the last inequality follows from Equation (K.21). We can pick $p_{k}$ and $p_{k^{\prime}}$ so that
@@
it then follows from Equation (K.22) that
@@
Combining the above display with Equation (K.21) finishes the proof. ∎
K.1.5 Proof of Lemma K.5
@@
For any $\mathbf{o}\in\widetilde{B}_{j}$, it follows from Equation (K.10) that
$f_{\bm{\omega}}(\mathbf{o})=\omega_{j}\varphi_{j}(\mathbf{o})\geq C_{\phi}M^{-\beta}\delta M^{\beta}=\delta C_{\phi}.$ (K.23)
For any $\bm{\omega}\in\Omega_{m}$ and $\delta>0$, combining Equation (K.23) with $\mathcal{R}_{T}^{\texttt{non}}$ as defined in Equation (K.6) yields that
$\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f_{\bm{\omega}})$ $=\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq\mathop{\mathrm{sign}}\left(f_{\bm{\omega}}(\bm{O}_{t})\right),V_{t}=0\right\}\left|f_{\bm{\omega}}(\bm{O}_{t})\right|\mid\mathcal{H}_{t-1}\right]$ (K.24)
$=\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq\mathop{\mathrm{sign}}\left(f_{\bm{\omega}}(\bm{O}_{t})\right),V_{t}=0\right\}\mathds{1}\left\{\bm{O}_{t}\in B_{j}\right\}\left|f_{\bm{\omega}}(\bm{O}_{t})\right|\mid\mathcal{H}_{t-1}\right]$
$\geq C_{\phi}\delta\sum_{j=1}^{m}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq\omega_{j},\bm{O}_{t}\in\widetilde{B}_{j},V_{t}=0\right\}\mid\mathcal{H}_{t-1}\right].$
@@
$\mathbb{P}\left(\bm{O}\in\widetilde{B}_{1}\right)$ $=\mathbb{P}\left(\phi_{\beta}\left(2M[\bm{O}-\bm{b}_{1}]\right)\geq\delta M^{\beta},\bm{O}\in B_{1}\right)$ (K.25)
$=\int_{B_{1}}\mathds{1}\left(\phi_{\beta}(2M(\mathbf{o}-\bm{b}_{1}))\geq\delta M^{\beta}\right)d\mathbf{o}$
$=\int_{B_{1}}\mathds{1}\left\{(1-2M\|\mathbf{o}-\bm{b}_{1}\|_{\infty})^{\beta}\geq\delta M^{\beta}\right\}d\mathbf{o}$
$=\int_{\left[0,\frac{1}{M}\right]^{d_{\texttt{non}}}}\mathds{1}\left(\max_{l\in[d_{\texttt{non}}]}\left|o_{l}-\frac{1}{2M}\right|\leq\frac{1}{2M}-\frac{1}{2}\delta^{1/\beta}\right)d\mathbf{o}$
$=\int_{\left[0,\frac{1}{M}\right]^{d_{\texttt{non}}}}\mathds{1}\left(\mathbf{o}\in\left[\frac{1}{2}\delta^{1/\beta},\frac{1}{M}-\frac{1}{2}\delta^{1/\beta}\right]^{d_{\texttt{non}}}\right)d\mathbf{o}$
@@
To handle $\mathcal{K}^{(\texttt{non})}$ as defined in Equation (K.1), take $\bm{\omega}$ and $\bm{\omega}^{\prime}$ so that they only differ in $\omega_{j}$, we have
@@
$\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[(f_{\bm{\omega}}(\bm{O}_{t})-f_{\bm{\omega}^{\prime}}(\bm{O}_{t}))^{2}\mathds{1}\left\{A_{t}=1,V_{t}=0\right\}\mid\mathcal{H}_{t-1}\right]$ (K.26)
$~~=4\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\varphi_{j}(\bm{O}_{t})^{2}\mathds{1}\left\{A_{t}=1,\bm{O}_{t}\in B_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
$~~\leq\frac{4C^{2}_{\phi}\delta^{2}}{M^{d_{\texttt{non}}}}+4\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\varphi_{j}(\bm{O}_{t})^{2}\mathds{1}\left\{A_{t}=1,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
$~~\leq\frac{4C^{2}_{\phi}\delta^{2}}{M^{d_{\texttt{non}}}}+4C_{\phi}^{2}M^{-2\beta}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}=1,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
where the first equality follows from the fact that $\bm{O}_{t}\in B_{j}$ already implies $V_{t}=0$. Similarly, for any $i\in[N]$, applying the above argument with Equation (K.25) to the pretrained data yields
$\mathbb{E}_{f_{\bm{\omega}}}\left[\left(f_{\bm{\omega}}\left(\bm{O}_{i}^{(0)}\right)-f_{\bm{\omega}^{\prime}}\left(\bm{O}_{i}^{(0)}\right)\right)^{2}\right]$ $\leq\frac{4C^{2}_{\phi}\delta^{2}}{M^{d_{\texttt{non}}}}+4C^{2}_{\phi}M^{-2\beta}\left(\frac{1}{M}-\delta^{1/\beta}\right)^{d_{\texttt{non}}}.$ (K.27)